# Highly Efficient Self-Adaptive Reward Shaping for Reinforcement Learning

### v3

@@ Abstract
Reward shaping is a technique in reinforcement learning that addresses the sparse-reward problem by providing more frequent and informative rewards. We introduce a self-adaptive and highly efficient reward shaping mechanism that incorporates success rates derived from historical experiences as shaped rewards. The success rates are sampled from Beta distributions, which dynamically evolve from uncertain to reliable values as data accumulates. Initially, the shaped rewards exhibit more randomness to encourage exploration, while over time, the increasing certainty enhances exploitation, naturally balancing exploration and exploitation. Our approach employs Kernel Density Estimation (KDE) combined with Random Fourier Features (RFF) to derive the Beta distributions, providing a computationally efficient, non-parametric, and learning-free solution for high-dimensional continuous state spaces. Our method is validated on various tasks with extremely sparse rewards, demonstrating notable improvements in sample efficiency and convergence stability over relevant baselines.
@@ 1 Introduction
Environments with extremely sparse rewards present notable challenges for reinforcement learning (RL). In such contexts, as the reward model lacks immediate signals, agents receive feedback only after long horizons, making the ability to quickly discover beneficial samples crucial for successful learning (Ladosz et al.,, 2022). To address this, a straightforward solution is to reconstruct the reward models by introducing auxiliary signals to assess the agent’s behavior, which has led to the popular technique of Reward Shaping (RS) (Strehl and Littman,, 2008; Gupta et al.,, 2022). Inverse reinforcement learning (IRL), which extracts reward functions from human knowledge or expert demonstrations, represents an intuitive approach within this framework Arora and Doshi, (2021). However, IRL heavily relies on extensive human input, which can be difficult to obtain especially in complex environments. Alternatively, fully autonomous approaches have emerged as an attractive direction.
@@
Automatically maintained reward shaping can be broadly categorized into two branches: intrinsic motivation-based rewards, which are task-agnostic, and inherent value-based rewards, which are typically task-specific. The former mainly introduces exploration bonuses to encourage agents to explore a wider range of states, commonly by rewarding novel or infrequently visited states (Burda et al.,, 2018; Ostrovski et al.,, 2017; Tang et al.,, 2017; Bellemare et al.,, 2016). While these methods effectively enhance exploration, they tend to overlook the internal values of the states. This can lead to the "noisy TV" problem, where agents fixate on highly novel but meaningless regions, thus trapping them in suboptimal behaviors (Mavor-Parker et al.,, 2022). In contrast, the latter leverages high-level heuristics to guide agents in extracting meaningful values from learning experiences, which helps stabilize convergence. (Ma et al., 2024b, ; Memarian et al.,, 2021; Trott et al.,, 2019).
@@
To overcome the limitations of existing RS methods and combine the advantages of exploration-encouraged and inherent value-based rewards, this paper introduces a novel Self-Adaptive Success Rate based reward shaping mechanism (SASR). The success rate, defined as the ratio of a state’s presence in successful trajectories to its total occurrences, works as an auxiliary reward distilled from historical experience. This success rate assesses a state’s contribution toward successful task completion, which closely aligns with the agent’s original objectives, offering informative guidance for learning. Furthermore, to mitigate overconfidence caused by deterministic success rates, we adopt Beta distributions to model success rates from a probabilistic perspective. Beta distributions enable a self-adaptive evolution of confidence in approximating a state’s success rate, ensuring the system gradually converges to reliable rewards as more data is collected, while preventing premature certainty. To derive Beta distributions, we use kernel density estimation (KDE) enhanced by random Fourier features (RFF) to estimate success and failure counts, formulating a highly efficient approach. The main contributions of this paper are summarized as follows:
@@
  • We propose SASR, an autonomous reward-shaping mechanism without the requirement of domain knowledge. By extracting a metric from historical experiences that aligns with the agent’s optimization objectives, SASR effectively augments the environmental rewards.
@@
  • A novel self-adaptive mechanism is introduced to achieve an efficient exploration-exploitation balance. Initially, low-confidence Beta distributions provide stochastic rewards that encourage exploration by diversifying policy optimization. As more experience accumulates, high-confidence Beta distributions offer more reliable and precise rewards to enhance exploitation.
@@
  • To derive the Beta distributions in continuous state spaces, we implement KDE with RFF. This formulates a non-parametric and learning-free approach, thereby achieving remarkably low computational complexity.
@@
  • SASR is evaluated on various extremely sparse-reward and continuous control tasks, significantly outperforming several baselines in sample efficiency, learning speed, and convergence stability.
@@ 2 Related Work
Reward shaping (RS) methods can generally be categorized based on the source of learning: either from human knowledge or agent’s own experiences. Techniques that derive reward models from human knowledge, such as Inverse Reinforcement Learning (IRL) (Arora and Doshi,, 2021; Ramachandran and Amir,, 2007; Ziebart et al.,, 2008; Hadfield-Menell et al.,, 2016) and Inverse Optimal Control (IOC) (Schultheis et al.,, 2021), aim to extract reward or objective functions from expert demonstrations. Following this, transferring the learned reward models to new tasks has received considerable efforts (Bıyık et al.,, 2022; Wu et al.,, 2021; Ellis et al.,, 2021; Cheng et al.,, 2021; Adamczyk et al.,, 2023). However, these methods rely heavily on human-generated data and often struggle to adapt to out-of-distribution scenarios. Thus, our focus shifts toward autonomous self-learning approaches, which can be further divided into intrinsic motivation-based and inherent value-based rewards based on the nature of the rewards.
@@
Intrinsic motivation based RS explores general heuristics or task-agnostic metrics to encourage exploration. Potential-based algorithms define the shaped reward as $\gamma\Phi(s^{\prime})-\Phi(s)$, where $\Phi(\cdot)$ is a potential function. This ensures the reward cancels out in the Bellman equation, preserving the optimal policy (Devlin and Kudenko,, 2012; Asmuth et al.,, 2008; Wiewiora,, 2003). However, designing the potential function is highly dependent on the environmental dynamics, making it more applicable to model-based RL. More commonly, methods incorporate exploration bonuses to reward novel states (Devidze et al.,, 2022; Badia et al.,, 2020; Hong et al.,, 2018; Eysenbach et al.,, 2019). Representatively, count-based strategies track visitation counts and assign higher rewards to less frequently visited states (Lobel et al.,, 2023; Machado et al.,, 2020; Fox et al.,, 2018; Fu et al.,, 2017; Martin et al.,, 2017). In continuous spaces, state counting is challenging, so Tang et al., (2017) introduced a hash function to discretize the state space, Bellemare et al., (2016) proposed pseudo-counts based on recording probabilities, and Ostrovski et al., (2017) used PixelCNN Van den Oord et al., (2016). Additionally, random network distillation based methods measure state novelty by neural networks (Yang et al.,, 2024; Burda et al.,, 2018), curiosity-driven approaches reward agents for encountering surprising or unpredictable states (Sun et al.,, 2022; Burda et al.,, 2019; Pathak et al.,, 2017). Although intrinsic motivation has proven effective in enhancing exploration, only considering novelty while ignoring the inherent values of states may lead to suboptimal policies. A notable example illustrating this limitation is the "noisy TV" problem (Mavor-Parker et al.,, 2022), where agents may become attracted to irrelevant, dynamic stimuli that generate high novelty but provide no useful information.
@@
Inherent value based RS, on the other hand, focuses on task-related signals that highlight how states contribute to achieving higher rewards and their underlying significance. For instance, Trott et al., (2019) introduced additional rewards based on the distance between a state and the target; Stadie et al., (2020) derived informative reward structures using a Self-Tuning Network to optimize guidance; Memarian et al., (2021) captured the preferences among different trajectories by ranking them via a trained classifier; Zheng et al., (2018) minimized the KL-divergence between learned and original rewards to align their distributions. Mguni et al., (2023) used an auxiliary agent competing against the original agent in a Markov game; Instead, Ma et al., 2024b introduced ReLara, a collaborative framework where an assistant reward agent automatically generates rewards to guide the policy agent. Moreover, incorporating multiple agents or hierarchical structures to share and transfer knowledge through synchronized reward functions is another promising research direction (Gupta et al.,, 2023; Ma et al., 2024c, ; Ma et al., 2024a, ; Park et al.,, 2023; Ma et al.,, 2023; Yi et al.,, 2022; Hu et al.,, 2020; Raileanu and Rocktäschel,, 2020).
While intrinsic motivation rewards guide exploration, inherent value rewards support convergence, Ma et al., 2024b combined these two advances by introducing a scheme that evolves the shaped rewards from random perturbations into meaningful metrics. Injecting random rewards during early training relaxes the assumption that only novel states should be rewarded, suggesting instead that any state can be randomly rewarded. Building on this idea, our study considers the success rate as an informative value. We sample this metric from evolving Beta distributions that progressively increase in confidence, enabling a self-adaptive exploration-exploitation balance. Compared with ReLara, which relies on a black-box agent, our method draws information from historical data, effectively reducing the delay of reward convergence and enhancing explainability.
@@ 3 Preliminaries
Reinforcement Learning (RL) aims to train an agent to interact with an environment, which is commonly modeled as a Markov Decision Process (MDP). An MDP represented as $\langle S,A,T,R,\gamma\rangle$, involves four main components: $S$ is the state space, $A$ is the action space, $T:S\times A\times S\rightarrow[0,1]$ is the probability of transitioning from one state to another given a specific action, and $R:S\rightarrow\mathbb{R}$ is the reward model. The objective in RL is to learn a policy $\pi(a|s)$ that maximizes the expected cumulative rewards $G=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t})]$, where $\pi(a|s)$ indicates the likelihood of selecting action $a$ in state $s$, and $\gamma$ is the discount factor (Sutton and Barto,, 2018).
Beta Distribution is defined on the interval $[0,1]$, making it ideal for modeling proportions or probabilities. It is parameterized by $\alpha$ and $\beta$, which are interpreted as prior counts of successes and failures of a binary outcome. The probability density function of a Beta-distributed variable $X$ is:
@@
where $B(\alpha,\beta)$ is a normalization factor. The key characteristic of Beta distribution is its adaptability: as more data accumulates, the values of $\alpha$ and $\beta$ increase, the distribution’s shape becomes narrower, thereby increasing confidence in the estimated probabilities. This feature is particularly beneficial in adaptive online learning, aligning closely with our objective of balancing exploration and exploitation.
Kernel Density Estimation (KDE) is a non-parametric approach to approximate the probability density function of a random variable from data samples. Given $n$ data points $\{x_{i}\}_{i=1}^{n}$, KDE smooths these points to approximate the density function as follows:
@@
where $h$ is the bandwidth, $K(\cdot)$ is a kernel function such as Gaussian kernel, Laplacian kernel, or Cauchy kernel. KDE is particularly useful in scenarios where the actual distribution is not well-defined or complex, such as continuous state spaces in our RL environments.
@@ 4 Methodology
We propose a Self-Adaptive Success Rate based reward shaping mechanism (SASR) to accelerate RL algorithms learning in extremely-sparse-reward environments. Figure 1 illustrates the principles of the SASR mechanism: The diagram consists of two parts representing the early and late learning stages. As experiences accumulate with learning progresses, the Beta distributions modeling the success rates evolve from being more stochastic to more deterministic. This autonomous adaption closely aligns with the agent’s exploration-exploitation balance. Section 4.1 introduces how Beta distributions evolve and how shaped rewards are generated from them. Additionally, to achieve highly efficient computation, we leverage KDE and RFF to estimate success and failure counts, which are used to derive the corresponding Beta distributions, as detailed in Section 4.2. Lastly, Section 4.3 presents the integration of SASR into the RL agent and the overall algorithmic flow.
@@ 4.1 Self-Adaptive Success Rate Sampling
where $N_{S}(s_{i})$ and $N_{F}(s_{i})$ denote the counts of state $s_{i}$ appearing in successful and failed historical trajectories, respectively, and $f(\cdot)$ is a mapping function that adjusts the magnitude of rewards: $f:[0,1]\rightarrow[R_{\min},R_{\max}]$.
Given $N_{S}(s_{i})$ and $N_{F}(s_{i})$, directly using a deterministic success rate $N_{S}(s_{i})/(N_{S}(s_{i})+N_{F}(s_{i}))$ may lead to overconfidence in the estimation of the true value. To address this, we model the success rate estimation from a probabilistic perspective, inspired by the principles of Thompson sampling, an effective algorithm widely used to balance exploration and exploitation in multi-armed bandit problems (Thompson,, 1933; Agrawal and Goyal,, 2012). We model the success rate of each state as an approximated variable in a Beta distribution, setting the shape parameters $\alpha=N_{S}(s_{i})+1$ and $\beta=N_{F}(s_{i})+1$:
@@
  • Encourage Exploration. In the early learning phases, the counts $N_{S}(s_{i})$ and $N_{F}(s_{i})$ are relatively low, resulting in broader Beta distributions and greater variability in sampled values, which injects significant randomness into the shaped rewards. These random rewards optimize the agent in various directions with tiny steps, thus shifting the anchors from which the stochastic policy samples actions. The strategy effectively broadens the range of policies and actions, thus encouraging the exploration. (Supporting experimental results in Section 5.2.)
  • Enhance Exploitation. In the late learning phases, as the counts $N_{S}(s_{i})$ and $N_{F}(s_{i})$ increase, the Beta distribution gradually sharpens, concentrating generated rewards closer to the true success rate. The evolution to a certain reward signal with higher confidence highly supports the agent’s exploitation, facilitating more efficient convergence towards optimal policies.
  • Consistent Optimization. The peak of the Beta distribution, computed as $N_{S}(s_{i})/(N_{S}(s_{i})+N_{F}(s_{i}))$, directly equals to the success rate. Meanwhile, the expectation, $(N_{S}(s_{i})+1)/(N_{S}(s_{i})+N_{F}(s_{i})+2)$, closely approximates the success rate. This ensures that, despite stochastic, the overall reward remains consistent with policy optimization.

### v4

@@ Abstract
Reward shaping is a reinforcement learning technique that addresses the sparse-reward problem by providing frequent, informative feedback. We propose an efficient self-adaptive reward-shaping mechanism that uses success rates derived from historical experiences as shaped rewards. The success rates are sampled from Beta distributions, which evolve from uncertainty to reliability as data accumulates. Initially, shaped rewards are stochastic to encourage exploration, gradually becoming more certain to promote exploitation and maintain a natural balance between exploration and exploitation. We apply Kernel Density Estimation (KDE) with Random Fourier Features (RFF) to derive Beta distributions, providing a computationally efficient solution for continuous and high-dimensional state spaces. Our method, validated on tasks with extremely sparse rewards, improves sample efficiency and convergence stability over relevant baselines.
@@ 1 Introduction
Environments with extremely sparse rewards present notable challenges for reinforcement learning (RL). In such contexts, as the reward model lacks immediate signals, agents receive feedback only after long horizons, making the ability to quickly discover beneficial samples crucial for successful learning (Ladosz et al., 2022). To address this, a straightforward solution is to reconstruct the reward models by introducing auxiliary signals that assess the agent’s behavior, which has led to the popular technique of Reward Shaping (RS) (Strehl & Littman, 2008; Gupta et al., 2022). Inverse reinforcement learning, which extracts reward functions from human knowledge or expert demonstrations, represents an intuitive approach within this framework Arora & Doshi (2021). However, it heavily relies on extensive human input, which can be difficult to obtain, especially in complex environments. Alternatively, fully autonomous approaches have emerged as an attractive direction.
@@
Automatically maintained reward shaping can be broadly categorized into two branches: intrinsic motivation-based rewards, which are task-agnostic, and inherent value-based rewards, which are typically task-specific. The former mainly introduces exploration bonuses to encourage agents to explore a wider range of states, commonly by rewarding novel or infrequently visited states (Burda et al., 2018; Ostrovski et al., 2017; Tang et al., 2017; Bellemare et al., 2016). While these methods effectively enhance exploration, they tend to overlook the internal values of the states. This can lead to the “noisy TV” problem, where agents fixate on highly novel but meaningless regions, thus trapping them in suboptimal behaviors (Mavor-Parker et al., 2022). In contrast, the latter leverages high-level heuristics to guide agents in extracting meaningful values from learning experiences, which helps stabilize convergence. However, these methods often struggle in early exploration due to non-directional guidance (Ma et al., 2024a; Memarian et al., 2021; Trott et al., 2019).
@@
To overcome the limitations of existing RS methods and combine the advantages of exploration-encouraged and inherent value-based rewards, this paper introduces a novel Self-Adaptive Success Rate-based reward shaping mechanism (SASR) [^1] The source code is accessible at: https://github.com/mahaozhe/SASR. The success rate, defined as the ratio of a state’s presence in successful trajectories to its total occurrences, serves as an auxiliary reward distilled from historical experience. This success rate assesses a state’s contribution toward successful task completion, which closely aligns with the agent’s original objectives, offering informative guidance for learning. Furthermore, to mitigate overconfidence caused by deterministic success rates, we adopt Beta distributions to model success rates from a probabilistic perspective. Beta distributions enable a self-adaptive evolution of confidence in approximating a state’s success rate, ensuring the system gradually converges to reliable rewards as more data is collected, while avoiding premature certainty. To derive Beta distributions, we use kernel density estimation (KDE) with random Fourier features (RFF) to efficiently estimate success and failure counts. The main contributions of this paper are summarized as follows:
@@
  • We propose SASR, an autonomous reward-shaping mechanism for sparse-reward environments. By deriving success rates from historical experiences aligned with the agent’s optimization objectives, SASR effectively augments the environmental rewards.
@@
  • We introduce a novel self-adaptive mechanism. Initially, low-confidence Beta distributions provide uncertain rewards, encouraging exploration by perturbing the reward function and assigning higher rewards to unvisited states. As more experience accumulates, high-confidence Beta distributions deliver more reliable and precise rewards to enhance exploitation.
@@
  • To derive Beta distributions in continuous state spaces, we use KDE with RFF, creating an efficient approach that eliminates the need for additional neural networks or models for learning the auxiliary shaped rewards, thereby achieving remarkably low computational complexity.
@@
  • SASR is evaluated on various extremely sparse-reward tasks, significantly outperforming several baselines in sample efficiency, learning speed, and convergence stability.
@@ 2 Related Work
Reward shaping (RS) methods can generally be categorized based on the source of learning: either from human knowledge or the agent’s own experiences. Techniques that derive reward models from human knowledge, such as Inverse Reinforcement Learning (IRL) (Arora & Doshi, 2021; Ramachandran & Amir, 2007; Ziebart et al., 2008; Hadfield-Menell et al., 2016) and Inverse Optimal Control (IOC) (Schultheis et al., 2021; Zhang et al., 2025), aim to extract reward or objective functions from expert demonstrations. Subsequently, transferring the learned reward models to new tasks has received considerable efforts (Bıyık et al., 2022; Wu et al., 2021; Ellis et al., 2021; Cheng et al., 2021; Adamczyk et al., 2023; Luo et al., 2024; Lyu et al., 2024). However, these methods rely heavily on human-generated data and often struggle adapting out-of-distribution scenarios. Thus, our focus shifts toward autonomous self-learning approaches, which can be further divided into intrinsic motivation-based and inherent value-based rewards depending on the nature of the rewards.
@@
Intrinsic motivation-based RS explores general heuristics or task-agnostic metrics to encourage exploration. Potential-based algorithms define the shaped reward as $\gamma\Phi(s^{\prime})-\Phi(s)$, where $\Phi(\cdot)$ is a potential function. This ensures the reward cancels out in the Bellman equation, preserving the optimal policy (Devlin & Kudenko, 2012; Asmuth et al., 2008; Wiewiora, 2003). However, designing the potential function highly depends on the environmental dynamics, making it more applicable to model-based RL. More commonly, methods incorporate exploration bonuses to reward novel states (Mahankali et al., 2024; Liu et al., 2025; Devidze et al., 2022; Badia et al., 2020; Hong et al., 2018; Eysenbach et al., 2019). Count-based strategies, for instance, track visitation counts and assign higher rewards to less frequently visited states (Lobel et al., 2023; Machado et al., 2020; Fox et al., 2018; Fu et al., 2017; Martin et al., 2017). In continuous spaces, state counting is challenging, so Tang et al. (2017) introduced a hash function to discretize the state space, Bellemare et al. (2016) proposed pseudo-counts based on recording probabilities, and Ostrovski et al. (2017) used PixelCNN Van den Oord et al. (2016) to simulate density. Additionally, random network distillation-based methods measure state novelty by neural networks (Yang et al., 2024b; Liu et al., 2024; Burda et al., 2018), while curiosity-driven approaches reward agents for encountering surprising or unpredictable states (Yang et al., 2024a; Sun et al., 2022; Burda et al., 2019; Pathak et al., 2017; Zhang et al., 2023). Although intrinsic motivation has proven effective in enhancing exploration, only considering novelty while ignoring the inherent values of states can lead to suboptimal policies.
@@
Inherent value based RS, on the other hand, focuses on task-related signals that highlight how states contribute to achieving higher rewards and their underlying significance. For instance, Trott et al. (2019) introduced additional rewards based on the distance between a state and the target; Stadie et al. (2020) derived informative reward structures using a Self-Tuning Network to optimize guidance; Memarian et al. (2021) captured the preferences among different trajectories by ranking them via a trained classifier; Zheng et al. (2018) minimized the KL-divergence between learned and original rewards to align their distributions. Mguni et al. (2023) used an auxiliary agent competing against the original agent in a Markov game; Ma et al. (2024a) introduced ReLara, a collaborative framework where a reward agent automatically generates rewards to guide the policy agent. Moreover, incorporating multiple agents or hierarchical structures to share and transfer knowledge through synchronized reward functions is another promising research direction (Park et al., 2023; Ma et al., 2024b; Gupta et al., 2023; Hu et al., 2020; Raileanu & Rocktäschel, 2020; Yi et al., 2022).
@@ 3 Preliminaries
Reinforcement Learning (RL) aims to train an agent to interact with an environment, which is commonly modeled as a Markov Decision Process (MDP). An MDP represented as $\langle S,A,T,R,\gamma\rangle$, involves four main components: $S$ is the state space, $A$ is the action space, $T:S\times A\times S\rightarrow[0,1]$ is the probability of transitioning from one state to another given a specific action, and $R:S\rightarrow\mathbb{R}$ is the reward model. The objective in RL is to learn a policy $\pi(a|s)$ that maximizes the expected cumulative rewards $G=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t})]$, where $\pi(a|s)$ indicates the likelihood of selecting action $a$ in state $s$, and $\gamma$ is the discount factor (Sutton & Barto, 2018).
Beta Distribution is defined on the interval $[0,1]$, making it ideal for modeling proportions or probabilities. It is parameterized by $\alpha$ and $\beta$, which represent prior counts of successes and failures of a binary outcome. The probability density function of a Beta-distributed variable $X$ is:
@@
where $B(\alpha,\beta)$ is the beta function. The key attribute of Beta distribution is its adaptability: as more data accumulates, the values of $\alpha$ and $\beta$ increase, narrowing the distribution’s shape and increasing confidence in the estimated probabilities. This feature is particularly useful in adaptive online learning, aligning with our objective of balancing exploration and exploitation.
Kernel Density Estimation (KDE) is a non-parametric method for approximating the probability density function of a random variable from data samples. Given $n$ data points $\{x_{i}\}_{i=1}^{n}$, KDE smooths these points to approximate the density function as follows:
@@
where $h$ is the bandwidth, and $K(\cdot)$ is a kernel function such as Gaussian kernel, Laplacian kernel, or Cauchy kernel. KDE is particularly useful in scenarios where the actual distribution is complex or poorly defined, such as in continuous state spaces in RL environments.
@@ 4 Methodology
We propose a Self-Adaptive Success Rate based reward shaping mechanism (SASR) to accelerate RL algorithms in extremely-sparse-reward environments. Figure 1 illustrates the principles of the SASR mechanism: The diagram consists of two parts representing the early and late learning stages. As experiences accumulate with learning progresses, the Beta distributions modeling the success rates evolve from being stochastic to deterministic. This autonomous adaption closely aligns with the agent’s exploration-exploitation balance. Section 4.1 introduces how Beta distributions evolve and how shaped rewards are generated from them. Additionally, to achieve highly efficient computation, we use KDE and RFF to estimate success and failure counts, which are used to derive the corresponding Beta distributions, as detailed in Section 4.2. Lastly, Section 4.3 presents the integration of SASR into the RL agent and the overall algorithmic flow.
@@ 4.1 Self-Adaptive Success Rate Sampling
where $N_{S}(s_{i})$ and $N_{F}(s_{i})$ denote the counts of state $s_{i}$ appearing in successful and failed historical trajectories, respectively. To enhance scalability and adaptability, $f(\cdot)$ is a linear scaling function that maps the original success rate from $[0,1]$ to a desired scale $[R_{min},R_{max}]$, making the magnitude of the shaped rewards more flexible, i.e., $f(x)=R_{min}+x\cdot(R_{max}-R_{min})$.
Given $N_{S}(s_{i})$ and $N_{F}(s_{i})$, directly using a deterministic success rate may lead to overconfidence in the estimation of the true value. To address this, inspired by the principles of Thompson sampling (Thompson, 1933; Agrawal & Goyal, 2012), we adopt a probabilistic perspective for success rate estimation. Specifically, the success rate of each state is approximated as a variable in a Beta distribution, with shape parameters set as $\alpha=N_{S}(s_{i})+1$ and $\beta=N_{F}(s_{i})+1$:
@@
  • Encourage Exploration. In the early phases, lower counts of $N_{S}(s_{i})$ and $N_{F}(s_{i})$ result in higher-variance Beta distributions, making the sampled rewards more stochastic. This acts as a noisy perturbation of the reward function. Given that the environmental rewards are mostly zero, this perturbation optimizes the agent in diverse directions through small adjustments, shifting the anchors from which the stochastic policy samples actions. Meanwhile, early-visited states are likely to fail, leading to a decrease in their success rates, while unvisited states retain the initial Beta distribution $Beta(1,1)$, receiving relatively higher rewards. This mechanism drives the agent to explore novel regions, aligning with the principles of intrinsic motivation.
  • Enhance Exploitation. In the later phases, as the counts $N_{S}(s_{i})$ and $N_{F}(s_{i})$ increase, the Beta distribution gradually sharpens, concentrating generated rewards around the true success rate. The more certain reward signals with higher confidence highly support the agent’s exploitation, facilitating faster convergence towards optimal policies.
  • Consistent Optimization. The peak of the Beta distribution, given by $N_{S}(s_{i})/(N_{S}(s_{i})+N_{F}(s_{i}))$, equals the success rate. Meanwhile, the expectation, $(N_{S}(s_{i})+1)/(N_{S}(s_{i})+N_{F}(s_{i})+2)$, closely approximates the success rate. This ensures that, despite stochasticity, the overall reward remains consistent with policy optimization.